# A class of charged and charged-Taub-NUT metrics in the presence of a massless scalar field and some of their astrophysical aspects

### v1

@@ Abstract
We consider a class of exact solutions to the Einstein equations in the presence of a scalar field, recently introduced in [1, 2], and derive their generalized form with dyonic charges using Harrison transformations. For specific parameter values, this class of metrics includes the charged Fisher-Janis-Newman-Winicour (FJNW) and $\gamma$-metrics. We then investigate the motion of neutral particles in the background of these metrics and derive the corresponding effective potential.
Next, by applying Ehlers transformations, we introduce the NUT parameter into the Reissner–Nordström metric in the presence of the scalar field. We also examine gravitational lensing, focusing on the effects of dyonic and NUT charges, as well as the scalar field, on the deflection angle of light. Finally, we explore the quasi-normal modes associated with this class of metrics.
1 Introduction
@@
The nature of a singularity is determined by the properties of the collapsing matter and the initial conditions of the system. For example, the gravitational collapse of a perfectly spherical, charged body leads to the formation of a Reissner–Nordström (RN) black hole—the simplest known example of a charged black hole. The influence of electric charge also extends to other spacetime geometries. One notable case is the $\gamma$-metric [40, 41, 42, 43], which describes a static, axially symmetric spacetime and has been extensively studied, including in its higher-dimensional generalizations [44]. Furthermore, a charged version of the $\gamma$-metric has been explored in [45].
@@
Recently, the $\gamma$-metric has been proposed as a viable model for describing the compact object at the center of our galaxy, as it successfully reproduces many of the observed features of this region [46, 47]. Another significant spacetime is the Fisher-Janis- Newman-Winicour (FJNW) metric [48, 49, 50], which represents a static, axisymmetric solution to the Einstein equations in the presence of a massless scalar field. The charged extension of the FJNW metric was first derived in [51], while correct form of its rotating counterpart was introduced in [2]. Many studies have examined the FJNW spacetime and the Einstein field equations minimally coupled to a scalar field [52, 53, 54, 1, 55, 2, 56, 36, 37, 39, 38].
@@
In [57, 58], Ernst reformulated the Einstein–Maxwell theory by introducing complex potential functions, leading to a more compact and symmetric representation of the field equations. These equations exhibit a rich symmetry structure and remain invariant under certain solution-generating techniques, notably the Harrison [59] and Ehlers [60] transformations. The Harrison transformation introduces a dyonic charge into the spacetime, while the Ehlers transformation incorporates the NUT (Newman–Unti–Tamburino) parameter. These transformations can also be extended to the Einstein theory coupled to a scalar field, as explored in [61, 62, 63].
@@
In this paper, we investigate a class of exact solutions to the Einstein equations recently proposed in [1, 2, 56]. By applying the Harrison and Ehlers transformations, we construct a new family of static, charged, Taub–NUT solutions that are asymptotically flat and include a massless scalar field. We then analyze the astrophysical properties of these newly derived charged Taub–NUT spacetimes.
@@
Geodesic motion plays a fundamental role in the study of compact astrophysical objects, offering crucial insights into the underlying geometry of spacetime. Among the various types of geodesics, circular orbits are of particular interest due to their relevance in observational astrophysics and their connection to stability analyses [64, 65]. In Sec. 2.2.1, we analyze the geodesic structure of the newly proposed class of metrics and identify the allowable range for one of the free parameters based on the conditions for physically viable circular motion.
@@
Another approach to understanding the properties of compact astrophysical objects is by studying their interactions with the surrounding environment. The perturbations of such objects give rise to the emission of gravitational waves, which can be analyzed through their corresponding quasi-normal modes (QNMs). QNMs are characterized by damped oscillations that evolve over time, providing valuable information about the object’s internal structure and the nature of its spacetime. For more details on QNMs and their calculation methods, refer to [88, 89, 90, 91, 92, 93, 94, 95, 96]. In Sec. 5, we employ the light ring method [97, 98, 99] to derive the QNMs.
The paper is organized as follows. In Sec. 2, we derive the charged versions of the metrics introduced in [1] using Harrison transformations. For specific choices of parameters, these metrics reduce to the charged Fisher–Janis–Newman–Winicour (FJNW) and charged $\gamma$-metrics. We then analyze geodesic motion to constrain the permissible range of the metric parameters. Additionally, we investigate the dynamics of neutral particles in the background of these spacetimes.
In Sec. 3, we apply both the Ehlers and Harrison transformations to construct a class of charged Taub-NUT metrics in the presence of a scalar field. Section 4 is devoted to the study of gravitational lensing, where we derive the deflection angle of light in scenarios with and without charge, the NUT parameter, and the scalar field. In Sec. 5, we compute the quasi-normal modes (QNMs) for specific parameter limits. Finally, Sec. 6 summarizes our findings and presents the conclusions.
In this article, we adopt natural units by setting $18\,\pi\,G=c=$ 1.
2 A class of static, charged metrics in the presence of a massless scalar field
@@
$S=\int d^{4}x\,\sqrt{-g}\,[R-\partial_{\mu}\varphi\,\partial^{\mu}\varphi-F_{\mu\nu}\,F^{\mu\nu}].$ (1)
@@
$\varphi(r)=\sqrt{\frac{1-\gamma^{2}-\nu}{2}}\;\ln\big{(}1-\frac{2\,m}{r}\big{)}.$ (8)
The metrics in Eq. (5) reduce to the FJNW metric for $\mu=1-\gamma$ and $\nu=0$, and to the $\gamma$-metric for $\mu=\gamma^{2}-\gamma$ and $\nu=1-\gamma^{2}.$ [^1] It has been recently pointed out in [55, 38] that the exact solutions that were originally obtained in [1] can also be derived by applying Buchdahl’s theorem [100, 101].
@@
Eriş and Gürses [61] have shown that, by adding a scalar field to the stationary and axisymmetric Einstein–electrovacuum solutions, the degree of integrability remains the same as in the vacuum case, and only the $\lambda$ function is modified in the LWP metric. Accordingly, the Ernst formulation remains fully applicable to Einstein’s equations with a minimally coupled scalar field, exactly as in the vacuum scenario [61, 62, 63].
@@
$\big{(}\text{Re}(\varepsilon)+\Phi\,\Phi^{\star}\big{)}\,\nabla^{2}\,\varepsilon=\nabla\,\varepsilon\,.\,(\nabla\,\varepsilon+2\,\Phi^{\star}\,\nabla\,\Phi),$ (13)
$\big{(}\text{Re}(\varepsilon)+\Phi\,\Phi^{\star}\big{)}\,\nabla^{2}\,\Phi=\nabla\,\Phi\,.\,(\nabla\,\varepsilon+2\,\Phi^{\star}\,\nabla\,\Phi).$
@@
$\varepsilon:=F-\lvert\Phi\rvert^{2}+i\,\chi,$ (14)
where $F$ is the metric coefficient given by (9), and the functions $\Phi$ and $\chi$ are defined as follows
$\Phi:=A_{t}+i\,\tilde{A}_{\phi},$ (15)
$\hat{\phi}\times\nabla\,\chi:=-\rho^{-1}\,F^{2}\,\nabla\,\omega-2\,\hat{\phi}\times Im(\Phi^{\star}\,\nabla\,\Phi),$
where $Im(\Phi^{\star}\,\nabla\,\Phi)$ represents the imaginary part of the function $\Phi^{\star}\,\nabla\,\Phi$ and $\hat{\phi}$ is the unit vector in the $\phi$ direction. Additionally, $A_{t}$ represents the first component of $A_{\mu}$ and $\tilde{A}_{\phi}$ satisfies the following equation
@@
where $A_{\phi}$ represents the fourth component of $A_{\mu}$. In [62, 63], Astorino demonstrated that, for the action in the presence of the scalar field, the scalar field decouples from Ernst’s equations, and an additional equation related to the scalar field, ($\Box\varphi(r)=0$), is added to Eqs. (13).
@@
$\partial_{\rho}\,\lambda=$ $\frac{\rho}{4\,(Re(\varepsilon)+\Phi\,\Phi^{*})^{2}}\,\big{[}(\partial_{\rho}\,\varepsilon+2\,\Phi^{*}\,\partial_{\rho}\,\Phi)\,(\partial_{\rho}\,\varepsilon^{*}+2\,\Phi\,\partial_{\rho}\,\Phi^{*})-(\partial_{z}\,\varepsilon+2\,\Phi^{*}\,\partial_{z}\,\Phi)\,(\partial_{z}\,\varepsilon^{*}$ (17)
$+2\,\Phi\,\partial_{z}\,\Phi^{*})\big{]}-\frac{\rho}{Re(\varepsilon)+\Phi\,\Phi^{*}}\,\big{[}\partial_{\rho}\,\Phi\,\partial_{\rho}\,\Phi^{*}-\partial_{z}\,\Phi\,\partial_{z}\,\Phi^{*}\big{]}+\frac{\rho}{2}\,\big{[}(\partial_{\rho}\,\varphi)^{2}-(\partial_{z}\,\varphi)^{2}\big{]},$
$\partial_{z}\,\lambda=$ $\frac{\rho}{4\,(Re(\varepsilon)+\Phi\,\Phi^{*})^{2}}\,\big{[}(\partial_{\rho}\,\varepsilon+2\,\Phi^{*}\,\partial_{\rho}\,\Phi)\,(\partial_{z}\,\varepsilon^{*}+2\,\Phi\,\partial_{z}\,\Phi^{*})+(\partial_{z}\,\varepsilon+2\,\Phi^{*}\,\partial_{z}\,\Phi)\,(\partial_{\rho}\,\varepsilon^{*}$ (18)
$+2\,\Phi\,\partial_{\rho}\,\Phi^{*})\big{]}-\frac{\rho}{Re(\varepsilon)+\Phi\,\Phi^{*}}\,\big{[}\partial_{\rho}\,\Phi\,\partial_{z}\,\Phi^{*}-\partial_{z}\,\Phi\,\partial_{\rho}\,\Phi^{*}\big{]}+\rho\,\big{[}\partial_{\rho}\,\varphi\,\partial_{z}\,\varphi\big{]},$
@@
$\varepsilon^{\prime}=\frac{\varepsilon}{1-2\,\alpha^{\star}\,\Phi-\lvert\alpha\rvert^{2}\,\varepsilon},\qquad\Phi^{\prime}=\frac{\alpha\,\varepsilon+\Phi}{1-2\,\alpha^{\star}\,\Phi-\lvert\alpha\rvert^{2}\,\varepsilon},$ (19)
@@
$\Phi^{\prime}=\frac{\alpha\,(r-2\,m)^{\gamma}}{r^{\gamma}-\lvert\alpha\rvert^{2}\,(r-2\,m)^{\gamma}}.$
@@
$F^{\prime}=\varepsilon^{\prime}+\lvert\Phi^{\prime}\rvert^{2}=\frac{\big{[}r\,(r-2\,m)\big{]}^{\gamma}}{\big{[}r^{\gamma}-(r-2\,m)^{\gamma}\;\lvert\alpha\rvert^{2}\big{]}^{2}},\qquad\chi^{\prime}=0.$ (22)
@@
×$\times\big{[}r\,(r-2\,m)+m^{2}\,\sin^{2}\theta\big{]}^{\nu}\,\big{\{}dr^{2}+[r\,(r-2\,m)]\,d\theta^{2}\big{\}}+\big{[}r^{\gamma}-(r-2\,m)^{\gamma}\;\lvert\alpha\rvert^{2}\big{]}^{2}$
×$\times\big{[}r\,(r-2\,m)\big{]}^{1-\gamma}\,\sin^{2}\theta\,d\phi^{2}.$
@@
$=\gamma\,q_{e}\,\int\,\frac{\mathcal{F}^{\gamma-1}}{\mathcal{R}^{2}\,r^{2}}.$
@@
$\varphi(r)=\sqrt{\frac{1-\gamma^{2}-\nu}{2}}\;\ln\big{(}\frac{r-m-\sqrt{m^{2}-q_{e}^{2}-q_{m}^{2}}}{r-m+\sqrt{m^{2}-q_{e}^{2}-q_{m}^{2}}}\big{)}.$ (35)
@@
The metrics presented in Eq. (31) constitute a class of Einstein–Maxwell solutions in the presence of a scalar field, which reduce to interesting known solutions in certain limits. If we set $\mu$ and $\nu$ as below
@@
metric (31) becomes the charged $\gamma$-metric. Now by choosing $\nu=0$ ($\mu=1-\gamma$), metric (31) represents the charged FJNW metric. When the electric and magnetic charges vanish ($q_{e}=q_{m}=0$), metric (31) reduces to metric (5), representing a spacetime with a scalar field, which is therefore not an electrovacuum solution of Einstein’s equations.
@@
$A^{\prime}=\frac{q}{r}\,dt,\quad\Rightarrow\quad\Phi^{\prime}=\frac{q}{r}.$ (40)
@@ 2.1 Singularities of the Metrics
Figure 1: The plot of the Ricci scalar in Eq. (41) as a function of $r$ is shown for $m=1$, $q_{e}=q_{m}=1/2$ and $\theta=\pi/6$. As is clear from the diagram, singularities occur at point $m+\sqrt{m^{2}-(q_{e}^{2}+q_{m}^{2})}$ and the Ricci scalar approaches zero at far distances.
2.2 Geodesics and Effective Potential
@@ 2.2.1 Geodesics
$\ddot{x}^{\alpha}+\Gamma_{\beta\,\sigma}^{\alpha}\,\dot{x}^{\beta}\,\dot{x}^{\sigma}=0,$ (44)
[…]
@@
2.2.2 Effective Potential of Neutral Particles
@@
$-\frac{\mathcal{F}^{\gamma}}{\mathcal{R}^{2}}\,\dot{t}^{2}+\mathcal{R}^{2}\,\mathcal{F}^{\mu-1}\,\mathcal{K}^{\nu}\,\dot{r}^{2}+\mathcal{R}^{2}\,\mathcal{F}^{\mu}\,\mathcal{K}^{\nu}\,r^{2}\,\dot{\theta}^{2}+\mathcal{R}^{2}\,\mathcal{F}^{1-\gamma}\,r^{2}\,\sin^{2}\theta\,\dot{\phi}^{2}=-\eta.$ (57)
@@
Figure 2: The plot of the effective potential in Eq. (63) as a function of $r$ is shown for $m=1$, $\eta=1$, $E^{2}/m_{0}^{2}=20$, and $L^{2}/m_{0}^{2}=0.1$. Plot (a) depicts graphs for the case where $\gamma\leq 1$. For $q_{e}=0.5$, $\gamma=0.6$, $\mu=0.4$, and $\nu=0$, the charged FJNW metric is represented by a solid red line. For $\gamma=0.6$, $\mu=-0.24$, and $\nu=0.64$, the charged $\gamma$-metric is shown by a black dashed line. For $\gamma=1$, $\mu=0$, and $\nu=0$, the RN metric is shown as a blue dot-dashed line. Plot (b) illustrates graphs for the case where $\gamma\geq 1$. For $q_{e}=0.5$, $\gamma=1.2$, $\mu=-0.2$, and $\nu=0$, the charged FJNW metric is shown by a solid red line. For $\gamma=1.2$, $\mu=0.24$, and $\nu=-0.44$, the charged $\gamma$-metric is shown by a black dashed line. For $\gamma=1$, $\mu=0$, and $\nu=0$, the RN metric is represented by a blue dot-dashed line. Plot (c) shows the case where $\gamma=0.8$, $\mu=0.2$, and $\nu=0$. For $q_{e}=0.99$, the charged FJNW metric is shown by a solid red line, and for $q_{e}=0$, the FJNW metric is represented by a black dashed line.
According to Fig. 2, in the case of $\gamma\leq 1$, the effective potential of neutral particles in the background of the charged $\gamma$-metric at distances close to the source is greater than the effective potential in the presence of RN and FJNW metrics. As expected, at far distances from the source, the effective potential reaches a constant value for all three metrics. In the case of $\gamma\geq 1$, unlike the previous case, the effective potential in the background of the charged $\gamma$-metric at distances close to the source is smaller than the effective potential in the presence of RN and FJNW metrics. Also, the presence of electric charge increases the effective potential near the gravitational source.
@@
3 A Class of Charged Taub-NUT Metrics in the Presence of a Scalar Field

### v2

@@ Abstract
Abstract
We consider a class of exact solutions to the Einstein equations in the presence of a scalar field, recently introduced in [1, 2], and derive their generalized form with dyonic charges using Harrison transformations. For specific parameter values, this class of metrics includes the charged Fisher-Janis-Newman-Winicour (FJNW) and Zipoy-Voorhees (ZV) metrics. We then investigate the motion of neutral particles in the background of these metrics and derive the corresponding effective potential.
Next, by applying Ehlers transformations, we introduce the NUT parameter into the Reissner-Nordström (RN) metric in the presence of the scalar field. We also examine gravitational lensing, focusing on the effects of dyonic and NUT charges, as well as the scalar field, on the deflection angle of light. Finally, we explore the quasi-normal modes associated with this class of metrics.
I Introduction
@@
The nature of a singularity is determined by the properties of the collapsing matter and the initial conditions of the system. For example, the gravitational collapse of a perfectly spherical, charged body leads to the formation of a Reissner-Nordström (RN) black hole—the simplest known example of a charged black hole. The influence of electric charge also extends to other spacetime geometries. One notable case is the Zipoy-Voorhees (ZV) metric [40, 41, 42, 43], which describes a static, axially symmetric spacetime and has been extensively studied, including in its higher-dimensional generalizations [44]. Furthermore, a charged version of the ZV metric ($\gamma$-metric) has been explored in [45].
@@
Recently, the ZV metric has been proposed as a viable model for describing the compact object at the center of our galaxy, as it successfully reproduces many of the observed features of this region [46, 47]. Another significant spacetime is the Fisher-Janis- Newman-Winicour (FJNW) metric [48, 49, 50], which represents a static, axisymmetric solution to the Einstein equations in the presence of a massless scalar field. The charged extension of the FJNW metric was first derived in [51]. Also, Bogush and Gal’tsov [52] demonstrated that previously proposed rotating FJNW solutions were incorrect, as they did not satisfy the Einstein field equations. They also proposed a rotating FJNW solution for specific values of the parameters and with a phantom scalar field. Later, in [2], for the first time a general physical rotating solution for FJNW has been found that is valid for all range of parameters with a real scalar field. Many studies have examined the FJNW spacetime and the Einstein field equations minimally coupled to a scalar field [53, 54, 55, 52, 1, 56, 2, 57, 36, 37, 39, 38].
@@
In [58, 59], Ernst reformulated the Einstein-Maxwell theory by introducing complex potential functions, leading to a more compact and symmetric representation of the field equations. These equations exhibit a rich symmetry structure and remain invariant under certain solution-generating techniques, notably the Harrison [60] and Ehlers [61] transformations. The Harrison transformation introduces a dyonic charge into the spacetime, while the Ehlers transformation incorporates the NUT parameter. These transformations can also be extended to the Einstein theory coupled to a scalar field, as explored in [62, 63, 64]. Recently, by using the Ernst scheme a new rotating solution for the vacuum Einstein’s equations was derived in [65].
@@
In this paper, we investigate a class of exact solutions to the Einstein equations recently proposed in [1, 2, 57]. By applying the Harrison and Ehlers transformations, we construct a new family of static, charged, Taub-NUT solutions that are asymptotically flat and include a massless scalar field. We then analyze the astrophysical properties of these newly derived charged Taub-NUT spacetimes.
@@
Geodesic motion plays a fundamental role in the study of compact astrophysical objects, offering crucial insights into the underlying geometry of spacetime. Among the various types of geodesics, circular orbits are of particular interest due to their relevance in observational astrophysics and their connection to stability analyses [66, 67]. In Sec. II.2.1, we analyze the geodesic structure of the newly proposed class of metrics and identify the allowable range for one of the free parameters based on the conditions for physically viable circular motion.
@@
Another approach to understanding the properties of compact astrophysical objects is by studying their interactions with the surrounding environment. The perturbations of such objects give rise to the emission of gravitational waves, which can be analyzed through their corresponding quasi-normal modes (QNMs). QNMs are characterized by damped oscillations that evolve over time, providing valuable information about the object’s internal structure and the nature of its spacetime. For more details on QNMs and their calculation methods, refer to [90, 91, 92, 93, 94, 95, 96, 97, 98]. In Sec. V, we employ the light ring method [99, 100, 101] to derive the QNMs.
The paper is organized as follows. In Sec. II, we derive the charged versions of the metrics introduced in [1] using Harrison transformations. For specific choices of parameters, these metrics reduce to the charged FJNW and charged ZV metrics. We then analyze geodesic motion to constrain the permissible range of the metric parameters. Additionally, we investigate the dynamics of neutral particles in the background of these spacetimes.
In Sec. III, we apply both the Ehlers and Harrison transformations to construct a class of charged Taub-NUT metrics in the presence of a scalar field. Section IV is devoted to the study of gravitational lensing, where we derive the deflection angle of light in scenarios with and without charge, the NUT parameter, and the scalar field. In Sec. V, we compute the quasi-normal modes (QNMs) for specific parameter limits. Finally, Sec. VI summarizes our findings and presents the conclusions.
In this article, we adopt natural units by setting $8\,\pi\,G=c=1$.
II A class of static, charged metrics in the presence of a massless scalar field
@@
$S=\intop\nolimits d^{4}x\,\sqrt{-g}\,[R-\partial_{\mu}\varphi\,\partial^{\mu}\varphi-F_{\mu\nu}\,F^{\mu\nu}].$ (1)
@@
[…]
The metrics in Eq. (5) reduce to the FJNW metric for $\mu=1-\gamma$ and $\nu=0$, and to the ZV metric for $\mu=\gamma^{2}-\gamma$ and $\nu=1-\gamma^{2}.$ [^1] It has been recently pointed out in [56, 38] that the exact solutions that were originally obtained in [1] can also be derived by applying Buchdahl’s theorem [102, 103].
@@
Eriş and Gürses [62] have shown that, by adding a scalar field to the stationary and axisymmetric Einstein-electrovacuum solutions, the degree of integrability remains the same as in the vacuum case, and only the $\lambda$ function is modified in the LWP metric. Accordingly, the Ernst formulation remains fully applicable to Einstein’s equations with a minimally coupled scalar field, exactly as in the vacuum scenario [62, 63, 64].
@@
[…]
@@
[…]
where $F$ is the metric coefficient given by (9), and the functions and $\chi$ are defined as follows
[…]
@@
where $A_{\phi}$ represents the fourth component of $A_{\mu}$. In [63, 64], Astorino demonstrated that, for the action in the presence of the scalar field, the scalar field decouples from Ernst’s equations, and an additional equation related to the scalar field, ($\square\varphi(r)=0$), is added to Eqs. (13).
@@
[…]
@@
[…]
@@
${}^{\prime}=\frac{\alpha\,(r-2\,m)^{\gamma}}{r^{\gamma}-\lvert\alpha\rvert^{2}\,(r-2\,m)^{\gamma}}.$
@@
$F^{\prime}=\varepsilon^{\prime}+\lvert{}^{\prime}\rvert^{2}=\frac{\big[r\,(r-2\,m)\big]^{\gamma}}{\big[r^{\gamma}-(r-2\,m)^{\gamma}\;\lvert\alpha\rvert^{2}\big]^{2}},\qquad\chi^{\prime}=0.$ (22)
@@
$\times\big[r\,(r-2\,m)+m^{2}\,\sin^{2}\theta\big]^{\nu}\,\big\{dr^{2}+[r\,(r-2\,m)]\,d\theta^{2}\big\}+\big[r^{\gamma}-(r-2\,m)^{\gamma}\;\lvert\alpha\rvert^{2}\big]^{2}$
$\times\big[r\,(r-2\,m)\big]^{1-\gamma}\,\sin^{2}\theta\,d\phi^{2}.$
@@
$=\gamma\,q_{e}\,\intop\nolimits\,\frac{\mathcal{F}^{\gamma-1}}{\mathcal{R}^{2}\,r^{2}}.$
@@
[…]
@@
The metrics presented in Eq. (31) constitute a class of Einstein-Maxwell solutions in the presence of a scalar field, which reduce to interesting known solutions in certain limits. If we set $\mu$ and $\nu$ as below
@@
metric (31) becomes the charged ZV metric. Now by choosing $\nu=0$ ($\mu=1-\gamma$), metric (31) represents the charged FJNW metric. When the electric and magnetic charges vanish ($q_{e}=q_{m}=0$), metric (31) reduces to metric (5), representing a spacetime with a scalar field, which is therefore not an electrovacuum solution of Einstein’s equations.
@@
$A^{\prime}=\frac{q}{r}\,dt,\quad\Rightarrow\quad{}^{\prime}=\frac{q}{r}.$ (40)
@@ II.1 Singularities of the Metrics
Figure 1: Plot of the Ricci scalar in Eq. (41) as a function of $r$, for $m=1$, $q_{e}=q_{m}=1/2$, and $\theta=\pi/6$. Singularities occur at $m+\sqrt{m^{2}-(q_{e}^{2}+q_{m}^{2})}$, and the Ricci scalar approaches zero at large $r$.
II.2 Geodesics and Effective Potential
@@ II.2.1 Geodesics
$\ddot{x}^{\alpha}+{}_{\beta\,\sigma}^{\alpha}\,\dot{x}^{\beta}\,\dot{x}^{\sigma}=0,$ (44)
$g_{\alpha\,\beta}\,\dot{x}^{\alpha}\,\dot{x}^{\beta}=-\eta,\qquad\eta=-1,\,0,\,1,$ (45)
@@
II.2.2 Effective Potential of Neutral Particles
@@
$\frac{\mathcal{F}^{\gamma}}{\mathcal{R}^{2}}\,\dot{t}^{2}+\mathcal{R}^{2}\,\mathcal{F}^{\mu-1}\,\mathcal{K}^{\nu}\,\dot{r}^{2}+\mathcal{R}^{2}\,\mathcal{F}^{\mu}\,\mathcal{K}^{\nu}\,r^{2}\,\dot{\theta}^{2}+\mathcal{R}^{2}\,\mathcal{F}^{1-\gamma}\,r^{2}\,\sin^{2}\theta\,\dot{\phi}^{2}=-\eta.$ (57)
@@
Figure 2: The plot of the effective potential in Eq. (63) as a function of $r$ is shown for $m=1$, $\eta=1$, $E^{2}/m_{0}^{2}=20$, and $L^{2}/m_{0}^{2}=0.1$. Plot (a) depicts graphs for the case where $\gamma\leq 1$. For $q_{e}=0.5$, $\gamma=0.6$, $\mu=0.4$, and $\nu=0$, the charged FJNW metric is represented by a solid red line. For $\gamma=0.6$, $\mu=-0.24$, and $\nu=0.64$, the charged ZV metric is shown by a black dashed line. For $\gamma=1$, $\mu=0$, and $\nu=0$, the RN metric is shown as a blue dot-dashed line. Plot (b) illustrates graphs for the case where $\gamma\geq 1$. For $q_{e}=0.5$, $\gamma=1.2$, $\mu=-0.2$, and $\nu=0$, the charged FJNW metric is shown by a solid red line. For $\gamma=1.2$, $\mu=0.24$, and $\nu=-0.44$, the charged ZV metric is shown by a black dashed line. For $\gamma=1$, $\mu=0$, and $\nu=0$, the RN metric is represented by a blue dot-dashed line. Plot (c) shows the case where $\gamma=0.8$, $\mu=0.2$, and $\nu=0$. For $q_{e}=0.99$, the charged FJNW metric is shown by a solid red line, and for $q_{e}=0$, the FJNW metric is represented by a black dashed line.
According to Fig. 2, in the case of $\gamma\leq 1$, the effective potential of neutral particles in the background of the charged ZV metric at distances close to the source is greater than the effective potential in the presence of RN and FJNW metrics. As expected, at far distances from the source, the effective potential reaches a constant value for all three metrics. In the case of $\gamma\geq 1$, unlike the previous case, the effective potential in the background of the charged ZV metric at distances close to the source is smaller than the effective potential in the presence of RN and FJNW metrics. Also, the presence of electric charge increases the effective potential near the gravitational source.
@@
III A Class of Charged Taub-NUT Metrics in the Presence of a Scalar Field